what is the x-intercept of y=x^2-2x-3?
Found 2 solutions by John10, ewatrrr:
Answer by John10(297)

When we find the x-intercept (let y =0) and vice versa
x^2 - 2x -3 =0
(x -3)(x + 1)= 0
x= 3 or x = -1
The x-intercept has the form (x,0): (3,0) and (-1,0)
It means that the equation cross the x-axis twice at x = 3 and x = -1

Answer by ewatrrr(24785)


Hi
x-intercepts (Note: when y = 0)
y=x^2-2x-3
x^2-2x-3=0
factoring
(x-3)(x+1)=0
(x-3)=0 x = 3
(x+1)=0 x = -1
x-intercepts: (3,0) and (-1,0)